OpenStudy (swaqqout_kid1):

A local company employs a varying number of employees each year, based on its needs. The labor costs for the company include a fixed cost of $30,383.00 each year, and $26,590.00 for each person employed for the year. For the next year, the company projects that labor costs will total $2,051,223.00. How many people does the company intend to employ next year? @sammixboo @sangya21

OpenStudy (anonymous):

As the labor charges are fixed for every year thus $2,051,223.00 - $30,383.00 = y no. of employees = y / $26,590.00= x

OpenStudy (anonymous):

Find x and y

sammixboo (sammixboo):

Well, we have \(\rm $30,383.00~+~$26,590.00p~=~$2,051,223.00\) We are finding the value of p, which is the people the company tend to employ next yea,

sammixboo (sammixboo):

To solve this, we must first subtract \(\rm $30,383.00\) on both sides of the equation to get p alone \(\rm $30,383.00~-~$30,383.00~+~$26,590.00p~=~$2,051,223.00~-~$30,383.00\) \(\rm $26,590.00p~=~$2,051,223.00~-~$30,383.00\) Now can you tell me what is \(\rm $2,051,223.00~-~$30,383.00\)?

OpenStudy (swaqqout_kid1):

2020840

sammixboo (sammixboo):

Yes, it equals \(\rm 2,020,840\), so now plug in \(\rm 2,020,840\) where \(\rm $2,051,223~-~$30,383\) was in the equation \(\rm 26,590p~=~ 2,020,840\) Now we must divide \(\rm 26,590\) on both sides of the equation to get p alone \(\rm \frac{26,590p}{26,590}~=~ \frac{2,020,840}{26,590}\) \(\rm p~=~ \frac{2,020,840}{26,590}\) Now, what is \(\rm 2,020,840\div 26,590\)

OpenStudy (swaqqout_kid1):

76

sammixboo (sammixboo):

Yes! So there would be around 76 people next year